# A Survey of Threats Against Voice Authentication and Anti-Spoofing Systems

### v2

@@ Abstract
journal: Computer Science Review
[…]
[a]organization=School of Information Technology, Deakin University, addressline=75 Pigdons Road, city=Waurn Ponds, postcode=3216, state=Victoria, country=Australia
1 Introduction
@@
This paper is organized as follows. Section 2 presents the motivation behind studying threats to voice authentication systems, emphasizing the growing reliance on voice-based security. Section 3 provides background on voice authentication systems and the typical threat models they face. Section 4 examines data poisoning attacks, discussing techniques that compromise training data to subvert model behavior. Section 5 delves into adversarial attacks, where carefully crafted audio inputs are designed to mislead voice authentication systems during inference. Section 6 investigates deepfake attacks, focusing on the creation of synthetic voices and their potential to impersonate legitimate users. Section 7 addresses adversarial spoofing attacks, which target anti-spoofing countermeasures to allow malicious audio to bypass security layers. Section 8 surveys prominent datasets, benchmarks, and competitions that support research and evaluation in this domain. Section 9 offers a comparison of the different attack strategies discussed. Section 10 outlines directions for future work, suggesting ways to enhance system robustness. Finally, Section 11 provides a conclusion, summarizing the key insights and implications of this review.
2 Motivation
@@
3 Background
3.1 List of Acronyms
To facilitate understanding, this subsection provides a list of acronyms commonly used throughout the document (see Table 1).
Table 1: List of Acronyms
@@ 3.1 List of Acronyms
3.2 Voice Authentication Systems Overview
@@ 3.2.1 System Architectures
3.2.2 Voice Authentication Tasks
@@
3.3 Attacks Definitions in Voice Authentication
@@
To better understand the evolving focus of the research community on these threats, Figure 2 shows the number of research studies published per year that specifically address attacks against voice authentication systems.
Figure 2: Number of research studies on attacks against voice authentication systems published per year.
@@
3.4 Attack Taxonomies
The landscape of attacks in voice authentication can be systematically categorized along several dimensions, as illustrated in Figure. 3.
@@ 3.4 Attack Taxonomies
3.5 Threat Models in Voice Authentication Attacks
@@
3.5.1 Data Poisoning Threat Models
@@
As discussed in Section 3.4, the attacker’s knowledge can vary: in a white-box setting, they may craft highly optimized poisoned inputs using full model access; in a gray-box scenario, poisoning is guided by partial system knowledge; and in a black-box setting, attacks rely on limited feedback or surrogate datasets. Poisoning goals may be targeted, such as embedding a backdoor to impersonate a specific user, or untargeted, aiming to degrade the model’s overall performance. Attack vectors include both clean-label and dirty-label strategies, depending on whether the injected samples are correctly or incorrectly labeled.
Further technical details and relevant literature will be discussed in Section 4.
3.5.2 Adversarial Threat Models
@@
According to the taxonomy in Section 3.4, adversarial attacks may occur under white-box conditions, where the attacker leverages full model access to compute gradients; gray-box settings, with partial system knowledge; or black-box scenarios, relying on queries or transferability. The objective may be targeted, forcing the system to accept an adversary as a specific user, or untargeted, simply inducing misclassification or denial of access. Effective adversarial examples maintain a trade-off between attack success and perceptual naturalness.
Further technical details and relevant literature will be discussed in Section 5.
3.5.3 Deepfake Voice Threat Models
@@
Further technical details and relevant literature will be discussed in Section 6.
3.5.4 Adversarial Spoofing Threat Models
@@
Further technical details and relevant literature will be discussed in Section 7.
4 Data Poisoning Attacks
Data poisoning attacks are a critical security concern for VAS, where adversaries manipulate training data to compromise a model’s integrity or availability [15, 16, 39]. With the increasing reliance on third-party data sources, including cloud-based services and web scraping, the risk of unintentionally incorporating poisoned data into training pipelines has increased significantly, particularly in the audio domain, where model training is computationally expensive and data-hungry [39, 15]. A summary of these attacks is illustrated in Table. 2.
4.1 Untargeted Data Poisoning Attacks
4.2 Targeted Data Poisoning Attacks
4.3 Backdoor Attacks
4.3.1 Trigger Mechanisms
4.3.2 Audio-Specific Triggers
4.3.3 Inaudible Triggers
Table 2: Summary of Data Poisoning and Backdoor Attacks on Voice Authentication Systems.
@@ 4.3.3 Inaudible Triggers
[42] 2022 Clean-label flip MFCC-based models $\geq$10% Less effective on ResNet/FFT, Kaldi shows robustness
@@
5 Adversarial Attacks
Adversarial attacks introduce subtle, often imperceptible, perturbations into speech signals to mislead Speaker Recognition Systems (SRS). Unlike spoofing attacks, which typically involve synthesized or replayed speech, adversarial attacks directly manipulate the input data to exploit and alter the model’s decision boundaries. As speaker recognition technology has transitioned from traditional methods, such as Gaussian Mixture Model-Universal Background Model (GMM-UBM) pipelines, to deep learning-based architectures like x-vectors [37] and ECAPA-TDNN [60], the sophistication of attack strategies has advanced in parallel, evolving from purely digital manipulations to practical over-the-air implementations. Table 3 summarizes key adversarial attacks on speaker recognition systems.
5.1 Foundational Work
5.2 White-Box Attacks
5.3 Black-Box Attacks
5.4 Physical Over-the-Air Attacks
PhyTalker [70] extended adversarial attacks to the real-time physical domain by injecting precisely timed, phoneme-level perturbations during live playback. PhyTalker achieved an 85% targeted ASR in over-the-air tests without requiring access to the target model’s parameters, demonstrating that physical attacks can compromise robust systems in real-time.
Table 3: Comparison of Key Adversarial Attacks on Speaker Recognition Systems.
@@ 5.4 Physical Over-the-Air Attacks
6 Deepfake Attacks
Deepfake speech attacks exploit advanced voice synthesis techniques, such as Voice Conversion (VC) and Text-to-Speech (TTS), to impersonate users in Speaker Recognition Systems (SRS). These attacks can bypass both automated authentication systems and human listeners by generating speech that closely mimics a target speaker’s voice, often using minimal training data. Unlike traditional replay spoofing, deepfake attacks can fabricate entirely new utterances, significantly expanding the threat landscape. As generative models become more powerful and accessible, these attacks pose a growing challenge to voice-based security. A summary of key deepfake attacks is presented in Table 3.
6.1 Foundational Work
6.2 Partial Fake (PF) Attacks
Table 4: Summary of Key Deepfake Attacks on Voice Authentication Systems.
@@ 6.2 Partial Fake (PF) Attacks
7 Adversarial Spoofing Attacks
Anti-spoofing countermeasures (CMs), also known as Presentation Attack Detection (PAD), are a critical defense layer in Speaker Verification (SV) pipelines, designed to distinguish bona fide speech from spoofed audio. However, recent studies have demonstrated that these systems are themselves highly vulnerable to adversarial attacks. These attacks introduce crafted, often imperceptible, perturbations to degrade the reliability of CMs, thereby increasing the false acceptance of synthetic or replayed speech. Table 5 presents a summary of these attacks.
7.1 White-Box Attacks
Early work by Liu et al. [98] applied white-box adversarial techniques, such as the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), against PAD models. Even minimal, imperceptible perturbations (i.e., small $\epsilon$) increased detector error rates to as high as 85–95%. These results revealed that, with access to model gradients, an attacker can nearly eliminate a PAD system’s ability to detect spoofed speech, effectively reducing its performance to that of random guessing.
7.2 Black-Box Attacks
Table 5: Summary of Adversarial Spoofing Attacks on Voice Authentication and Anti-Spoofing Systems.
@@ 7.2 Black-Box Attacks
8 Evaluation
8.1 Popular Datasets
Evaluating VAS and studying their vulnerabilities requires access to diverse and well-annotated datasets. Over the past decade, several benchmark datasets have been developed to support research in SV, spoofing detection, and adversarial robustness. An overview of dataset distribution per field is illustrated in Figure 5, offering a visual breakdown of the number of datasets categorized by their primary research purpose. Additionally, A detailed comparison of these datasets is provided in Table.  6.
Table 6: Comparison of Voice Authentication and Spoofing Datasets.
@@ 8.1 Popular Datasets
8.2 Evaluation Metrics
@@
8.3 Competitions
Table 7 provides a comparative overview of the major anti-spoofing competitions, summarizing their key characteristics, attack types, and contributions to the field.
8.3.1 Speaker Verification Competitions
8.3.2 Anti-spoofing Competitions
Table 7: Summary of Major Anti-Spoofing Competitions.
@@ 8.3.2 Anti-spoofing Competitions
9 Comparison and Limitations
This section provides a comparative analysis of the attack methods discussed throughout the previous sections. Table 8 summarizes the common limitations identified across all reviewed attack categories, highlighting shared vulnerabilities and constraints that hinder their practical deployment in real-world voice authentication systems.
9.1 Data Poisoning
@@
9.2 Adversarial Attacks
Among the studies reviewed in Section 5, adversarial attacks on SRSs largely focus on white-box and black-box scenarios. White-box methods leverage model gradients, while black-box methods aim for query-efficient, imperceptible attacks. Despite innovations such as psychoacoustic modeling and ensemble-based surrogates, practical limitations remain due to reliance on score access, high query budgets, or constrained physical-world setups.
@@ 9.2 Adversarial Attacks
9.3 Deepfake Attacks
As outlined in Section 6, deepfake-based attacks present an increasingly serious threat to SRSs, leveraging generative models for voice imitation or partial utterance replacement. Unlike traditional adversarial attacks that rely on optimization or model access, deepfake methods often succeed in fully black-box, zero-shot settings—making them highly practical and scalable. Early work using voice morphing and more recent partial spoofing (PF) techniques confirm the vulnerability of both machine and human verification systems, even with minimal attacker knowledge.
9.4 Adversarial Attacks against CMs
As discussed in Section 7, most reviewed defenses rely on passive, model-agnostic strategies, yet they fall short against modern, adaptive attack techniques. Recent adversarial methods, discussed in Section 5, exploit subtle, non-obvious audio features or leverage universal perturbations that evade traditional PAD systems—even under physical conditions. Notably, none of the defenses evaluated address real-time constraints or defense latency, leaving systems vulnerable to fast, transferable attacks.
Table 8: Limitations Comparison for Data Poisoning, Adversarial, Deepfake, and Adversarial Spoofing Attacks.
@@ 9.4 Adversarial Attacks against CMs
10 Future Work
@@
10.1 Data Poisoning Attacks
@@ 10.3 Deepfake Attacks
10.4 Adversarial Spoofing Attacks
@@ 11 Conclusion
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used ChatGPT 4.0 to check grammar, correct language, and paraphrase certain sentences to improve readability. After using this tool, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.
Data availability
No data was used for the research described in the article.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v4

@@ Abstract
I Introduction
@@
This paper is organized as follows. Section II presents the motivation behind studying threats to voice authentication systems, emphasizing the growing reliance on voice-based security. Section III provides background on voice authentication systems and the typical threat models they face. Section IV examines data poisoning attacks, discussing techniques that compromise training data to subvert model behavior. Section V delves into adversarial attacks, where carefully crafted audio inputs are designed to mislead voice authentication systems during inference. Section VI investigates deepfake attacks, focusing on the creation of synthetic voices and their potential to impersonate legitimate users. Section VII addresses adversarial spoofing attacks, which target anti-spoofing countermeasures to allow malicious audio to bypass security layers. Section VIII surveys prominent datasets, benchmarks, and competitions that support research and evaluation in this domain. Section IX offers a comparison of the different attack strategies discussed. Section X outlines directions for future work, suggesting ways to enhance system robustness. Finally, Section XI provides a conclusion, summarizing the key insights and implications of this review.
II Motivation
@@
III Background
III-A List of Acronyms
To facilitate understanding, this subsection provides a list of acronyms commonly used throughout the document (see Table I).
TABLE I: List of Acronyms
@@ III-A List of Acronyms
III-B Voice Authentication Systems Overview
@@ III-B1 System Architectures
III-B2 Voice Authentication Tasks
@@
III-C Attacks Definitions in Voice Authentication
@@
To better understand the evolving focus of the research community on these threats, Figure 2 shows the cumulative number of research studies published over time that specifically address attacks against voice authentication systems.
Figure 2: Cumulative number of research studies on attacks against voice authentication systems.
@@
III-D Attack Taxonomies
The landscape of attacks in voice authentication can be systematically categorized along several dimensions, as illustrated in Fig. 3.
@@ III-D Attack Taxonomies
III-E Threat Models in Voice Authentication Attacks
@@
III-E1 Data Poisoning Threat Models
@@
As discussed in Section III-D, the attacker’s knowledge can vary: in a white-box setting, they may craft highly optimized poisoned inputs using full model access; in a gray-box scenario, poisoning is guided by partial system knowledge; and in a black-box setting, attacks rely on limited feedback or surrogate datasets. Poisoning goals may be targeted, such as embedding a backdoor to impersonate a specific user, or untargeted, aiming to degrade the model’s overall performance. Attack vectors include both clean-label and dirty-label strategies, depending on whether the injected samples are correctly or incorrectly labeled.
Further technical details and relevant literature will be discussed in Section IV.
III-E2 Adversarial Threat Models
@@
According to the taxonomy in Section III-D, adversarial attacks may occur under white-box conditions, where the attacker leverages full model access to compute gradients; gray-box settings, with partial system knowledge; or black-box scenarios, relying on queries or transferability. The objective may be targeted, forcing the system to accept an adversary as a specific user, or untargeted, simply inducing misclassification or denial of access. Effective adversarial examples maintain a trade-off between attack success and perceptual naturalness.
Further technical details and relevant literature will be discussed in Section V.
III-E3 Deepfake Voice Threat Models
@@
Further technical details and relevant literature will be discussed in Section VI.
III-E4 Adversarial Spoofing Threat Models
@@
Further technical details and relevant literature will be discussed in Section VII.
IV Data Poisoning Attacks
TABLE II: Summary of Data Poisoning and Backdoor Attacks against Voice Authentication Systems
@@ IV Data Poisoning Attacks
[41] 2022 Clean-label flip MFCC-based models $\geq 10\%$ Less effective on ResNet/FFT, Kaldi shows robustness
@@
Data poisoning attacks are a critical security concern for VAS, where adversaries manipulate training data to compromise a model’s integrity or availability [15, 16, 53]. With the increasing reliance on third-party data sources, including cloud-based services and web scraping, the risk of unintentionally incorporating poisoned data into training pipelines has increased significantly, particularly in the audio domain, where model training is computationally expensive and data-hungry [53, 15]. A summary of these attacks is illustrated in Table. II.
IV-A Untargeted Data Poisoning Attacks
IV-B Targeted Data Poisoning Attacks
IV-C Backdoor Attacks
IV-C1 Trigger Mechanisms
IV-C2 Audio-Specific Triggers
IV-C3 Inaudible Triggers
V Adversarial Attacks
TABLE III: Comparison of Key Adversarial Attacks on Speaker Recognition Systems.
@@ V Adversarial Attacks
Adversarial attacks introduce subtle, often imperceptible, perturbations into speech signals to mislead Speaker Recognition Systems (SRS). Unlike spoofing attacks, which typically involve synthesized or replayed speech, adversarial attacks directly manipulate the input data to exploit and alter the model’s decision boundaries. As speaker recognition technology has transitioned from traditional methods, such as Gaussian Mixture Model-Universal Background Model (GMM-UBM) pipelines, to deep learning-based architectures like x-vectors [37] and ECAPA-TDNN [69], the sophistication of attack strategies has advanced in parallel, evolving from purely digital manipulations to practical over-the-air implementations. Table III summarizes key adversarial attacks on speaker recognition systems.
V-A Foundational Work
V-B White-Box Attacks
V-C Black-Box Attacks
V-D Physical Over-the-Air Attacks
PhyTalker [68] extended adversarial attacks to the real-time physical domain by injecting precisely timed, phoneme-level perturbations during live playback. PhyTalker achieved an 85% targeted ASR in over-the-air t
VI Deepfake Attacks
TABLE IV: Summary of Key Deepfake Attacks on Voice Authentication Systems.
@@ VI Deepfake Attacks
Deepfake speech attacks exploit advanced voice synthesis techniques, such as Voice Conversion (VC) and Text-to-Speech (TTS), to impersonate users in Speaker Recognition Systems (SRS). These attacks can bypass both automated authentication systems and human listeners by generating speech that closely mimics a target speaker’s voice, often using minimal training data. Unlike traditional replay spoofing, deepfake attacks can fabricate entirely new utterances, significantly expanding the threat landscape. As generative models become more powerful and accessible, these attacks pose a growing challenge to voice-based security. A summary of key deepfake attacks is presented in Table III.
VI-A Foundational Work
VI-B Partial Fake (PF) Attacks
VII Adversarial Spoofing Attacks
TABLE V: Summary of Adversarial Spoofing Attacks on Voice Authentication and Anti-Spoofing Systems.
@@ VII Adversarial Spoofing Attacks
Anti-spoofing countermeasures (CMs), also known as Presentation Attack Detection (PAD), are a critical defense layer in Speaker Verification (SV) pipelines, designed to distinguish bona fide speech from spoofed audio. However, recent studies have demonstrated that these systems are themselves highly vulnerable to adversarial attacks. These attacks introduce crafted, often imperceptible, perturbations to degrade the reliability of CMs, thereby increasing the false acceptance of synthetic or replayed speech. Table V presents a summary of these attacks.
VII-A White-Box Attacks
Early work by Liu et al. [98] applied white-box adversarial techniques, such as the FGSM and PGD, against PAD models. Even minimal, imperceptible perturbations (i.e., small $\epsilon$) increased detector error rates to as high as 85–95%. These results revealed that, with access to model gradients, an attacker can nearly eliminate a PAD system’s ability to detect spoofed speech, effectively reducing its performance to that of random guessing.
VII-B Black-Box Attacks
VIII Evaluation
TABLE VI: Comparison of Voice Authentication and Spoofing Datasets.
@@ VIII Evaluation
VIII-A Popular Datasets
Evaluating VAS and studying their vulnerabilities requires access to diverse and well-annotated datasets. Over the past decade, several benchmark datasets have been developed to support research in SV, spoofing detection, and adversarial robustness. An overview of dataset distribution per field is illustrated in Figure 5, offering a visual breakdown of the number of datasets categorized by their primary research purpose. Additionally, A detailed comparison of these datasets is provided in Table.  VI.
VIII-B Evaluation Metrics
@@
VIII-C Competitions
Table VII provides a comparative overview of the major anti-spoofing competitions, summarizing their key characteristics, attack types, and contributions to the field.
TABLE VII: Summary of Major Anti-Spoofing Competitions
@@ VIII-C Competitions
VIII-C1 Speaker Verification Competitions
VIII-C2 Anti-spoofing Competitions
IX Comparison and Limitations
This section provides a comparative analysis of the attack methods discussed throughout the previous sections. Table VIII summarizes the common limitations identified across all reviewed attack categories, highlighting shared vulnerabilities and constraints that hinder their practical deployment in real-world voice authentication systems.
IX-A Data Poisoning
@@
IX-B Adversarial Attacks
Among the studies reviewed in Section V, adversarial attacks on SRSs largely focus on white-box and black-box scenarios. White-box methods leverage model gradients, while black-box methods aim for query-efficient, imperceptible attacks. Despite innovations such as psychoacoustic modeling and ensemble-based surrogates, practical limitations remain due to reliance on score access, high query budgets, or constrained physical-world setups.
@@ IX-B Adversarial Attacks
TABLE VIII: Limitations Comparison for Data Poisoning, Adversarial, Deepfake, and Adversarial Spoofing Attacks
@@
IX-C Deepfake Attacks
As outlined in Section VI, deepfake-based attacks present an increasingly serious threat to SRSs, leveraging generative models for voice imitation or partial utterance replacement. Unlike traditional adversarial attacks that rely on optimization or model access, deepfake methods often succeed in fully black-box, zero-shot settings—making them highly practical and scalable. Early work using voice morphing and more recent partial spoofing (PF) techniques confirm the vulnerability of both machine and human verification systems, even with minimal attacker knowledge.
IX-D Adversarial Spoofing Attacks
As discussed in Section VII, most reviewed defenses rely on passive, model-agnostic strategies, yet they fall short against modern, adaptive attack techniques. Recent adversarial methods, discussed in Section V, exploit subtle, non-obvious audio features or leverage universal perturbations that evade traditional PAD systems—even under physical conditions. Notably, none of the defenses evaluated address real-time constraints or defense latency, leaving systems vulnerable to fast, transferable attacks.
X Future Work
@@
X-A Data Poisoning Attacks
@@ X-C Deepfake Attacks
X-D Adversarial Spoofing Attacks